The difference between two positive integers is 36. One integer is three times as great as the other. Find the integers
Answer by checkley77(12844)

x-y=36
x=3y
3y-y=36
2y=36
y=36/2
y=18 ans. for the smaller integer.
x=3*18
x=54
Proof:
54-18=36
36=36